OpenStudy (anonymous):

what is the domain of the function: f(x) = 6/(x-4)

OpenStudy (ash2326):

we have f(x)= 6/(x-4) x can't be 4 that'll make f(x) infinity so domain is all real numbers except 4 R-{4}

OpenStudy (anonymous):

technically not infinity, it'll be undefined

OpenStudy (anonymous):

how do i state it correctly?

OpenStudy (anonymous):

\[D=\left\{ x \neq4, x \in R \right\}\]

OpenStudy (anonymous):

*D: {x≠4,x∈R}

OpenStudy (anonymous):

thanks, can I just put \[\neq\] 4 ?

OpenStudy (anonymous):

i mean x is not equal to 4

OpenStudy (anonymous):

depends on what your teacher wants

OpenStudy (anonymous):

well how do you do the work to see f(x) = 6/(x-4) please

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i still don't totally understand domain, please help a bit more

OpenStudy (anonymous):

domain is all the possible x values of a function which gives f(x) a value

OpenStudy (anonymous):

you cannot divide a number by 0, because it will be undefined, thus f(x) won't be defined at that point

OpenStudy (anonymous):

if x=4, f(x)=6/0 which is undefined

OpenStudy (anonymous):

*f(4)=6/0 which is undefined

OpenStudy (anonymous):

thank you, so I wrote that the answer is undefined

OpenStudy (anonymous):

does that mean 6 divided by 6-4?

OpenStudy (anonymous):

6/(4-4)

Directrix (directrix):

Take a look at, say, the first minute of this video. Note the two basic guidelines on what to watch for when finding domains. http://www.youtube.com/watch?v=sbvSBH2Mo20